# Unexpected but informative: What fixation-related potentials tell us about the processing of confusing program code

### v2

@@ 3 Discussion
The primary goal of this study was to apply an FRP approach – an established technique from psycholinguistics – to achieve a better understanding of the neurocognitive processes that occur when reading confusing program code (i.e., code that contains atoms of confusion) by means of exploring parallels between the processing of program code and natural language processing in an ecologically valid experimental setting. Therefore, we investigated whether atoms of confusion trigger electrophysiological responses, which turned out to be similar to those engaged by unexpected but plausible inputs during language comprehension. In particular, our FRP analysis revealed a late positivity with a frontal scalp topography. It onsets around after the first eye fixation in the area containing the atom of confusion and was significantly larger for confusing code compared to clean code snippets. Notably, this component resembles other late frontal positivities found in natural language processing studies, in which the late frontal positivity is elicited by unexpected words, often but not necessarily in highly constraining contexts [45, 46, 47]. As outlined in Section 1, the late frontal positivity can be dissociated from the P600, a late ERP component, that is elicited by words that are perceived as highly implausible during online comprehension [27, 48].
@@
So what does the late frontal positivity reveal about the processes involved in the processing of an atom of confusion? In the literature on natural language comprehension, the predominant account for a late frontal positivity holds that it reflects expectancy mismatch related processing [45, 46] or late processing costs, due to the perceived difference between a prediction and the bottom-up input in highly constraining sentences [49, 50]. However, a range of studies also observed that unexpected words within weakly constraining sentences can elicit a late frontal positivity [51, 52, 26, 53, 54, 55]. Brothers et al. [34, 56] proposed that the essential properties of a stimulus that lead to a late frontal positivity might be the presence of an unexpected but plausible word, which causes a substantial update of the situation model. More specifically, Kuperberg et al. [57] suggest that late frontal positivities arise in situations in which there is first uncertainty about the situation model, which triggers retrieval of information from long term memory, which subsequently enables the comprehender to accumulate confidence about being able to successfully update their situation model. In the case of high contextual constraint and unexpected lexical material, this occurs because the unexpected word first reduces the confidence in the situation model, triggers retrieval from long-term memory, and finally leads to a high confidence decision that the situation model can be successfully updated [57]. In the case of a low-constraint context and unexpected but plausible lexical input, retrieval of relevant event-level information from long-term memory allows the comprehender to accumulate confidence about how to update their situation model. Kuperberg et al. [57] point out that the late frontal positivity effect in language processing seems to depend on the task, how engaged an individual is in deeply understanding the situation, and whether they can achieve confidence about the situation model based on the information they retrieved from long-term memory.
@@
These ideas align well with a plausible mechanism for processing the atoms of confusion: Atoms of confusion are infrequent in code [11] and hence generally unexpected, but they are not implausible, or incorrect, as they reflect syntactically correct code patterns that can be unambiguously executed by the computer. However, when a comprehender encounters an atom of confusion, they are likely to be unsure at first about its effect on the program execution and variables they need to keep track of for the task (i.e., they are unsure about how they need to update their situation model), which is in line with observations that relate confusing code patterns with increased cognitive load [16]. As a result, comprehenders access their long-term memory to remember how exactly to interpret the atom of confusion. If this retrieval is successful, the comprehender can confidently update their situation model. In the case of code without atom, the situation model will also be updated, but there is no point of uncertainty followed by confidence on how to do that. This account also generates new predictions to be tested in future studies: We would expect that the late frontal positivity should only be present in trials where the comprehender could successfully resolve the ambiguity. It should hence not show up in situations in which the atom of confusion contains a syntactically invalid code element, for which we would expect a P600 with a posterior scalp topography [42]. Also, we do not expect a late frontal positivity to show up when the comprehender remains uncertain about the semantics of the atom, or when they would be given a different task that does not require updating the situation model.
@@
In summary, we presented, to our knowledge, the first-ever study that employs an FRP approach to explore the processing of confusing elements in program code. The results suggest that the brain engages highly similar mechanisms for processing unexpected but plausible elements in program code and natural language comprehension. This implies that studies on program comprehension can rely on established ERP components in natural language processing. The FRP approach offers new research perspectives with high ecological validity, focusing on the temporal characteristics of online program code comprehension. These results are significant for software engineering research, as they provide a foundation for a better understanding of code comprehension. Based on this common methodological ground, methods and findings from program comprehension research may be transferred back to research on natural language comprehension. The findings of our study, in particular, and this line of research, in general, also have far-reaching implications for the practice of software engineering. By shedding light on how programmers’ brains process and respond to confusing code, we can enable programmers to develop and maintain more secure and reliable software. Specifically, our line of research supports programming practice by providing empirical evidence for coding conventions that avoid confusion-inducing code patterns. Moreover, programming language designers can leverage these insights to design consistent language constructs that minimize the potential for confusion and resulting severe mistakes. Finally, educators can incorporate these findings into their teaching practices, ultimately educating future programmers to apply to more effective software development practices.
@@ 4.1 Participants
To estimate the number of participants required to find the hypothesized effect, we performed a power analysis with , , and (based on the reported ) for the late frontal positivity detected by Thornhill and Peterson [26] for high-constraint unrelated sentence endings. This analysis estimated a minimum of participants to correctly measure an existing effect (two-sided t-test). Due to various uncertainties in our setup and for counterbalancing of the three version blocks across participants, we increased our sample to participants.
We recruited participants at our host institution (Saarland University) via e-mail lists, flyers, and (online) bulletin boards. We required participants to have, at least, an undergraduate degree in computer science or a related field. We further asked for basic experience with an object-oriented programming language (e.g., Java or C++), as the selected snippets use minimal language-specific operations. We only invited right-handed participants.
@@
Overall, we invited participants to our experiment. To ensure high data quality, we excluded two participants from our analysis due to hardware malfunctions during the experiment or frequent excessive artifacts in the EEG data (see Section 4.5.4 and 4.5.5). out of the included participants were female, and male, with a total age range from to years, and with participants being younger than . In Table 2, we provide an overview of our included participants’ demographics and their programming experience.
@@ 4.2 Experimental variables
The independent variable of our experiment is the snippet type, which has two conditions: either confusing or clean. As dependent variables, we measured for each snippet the comprehension time, the answer correctness, the subjective difficulty rating, the eye movements via an eye tracker, and the brain activity via an EEG system (see Figure 2).
@@ 4.3 Stimuli
The code snippets in our study are based on the snippet pairs from the study of Langhout and Aniche [13], which we adjusted to reduce the fixations and confounding neurocognitive processes in the FRP analysis. In a pilot study, we rated these snippet pairs considering three criteria regarding the snippet structure, area of the atom of confusion, and behavioral response (cf. supplementary information, Appendix A). Specifically, we required that, within each snippet pair, the snippets have a comparable structure to reduce the influence of confounding factors. This was determined by the number and length of the code lines. Furthermore, to facilitate the selection of suitable fixations for FRP analysis, each snippet in a pair must contain a single clear area of interest that contains either the atom of confusion (for confusing snippets) or a clarified equivalent (for clean snippets). Finally, we tested whether the snippets were of a comparative complexity and whether they were, in principle, understandable to students of our target population (familiarity with the syntax of Java) in a pilot study. To this end, we set a threshold of answer correctness or subjective difficulty rating as easy for either condition. Snippets that showed a clear violation of the expected behavior by less correctness or fewer easy ratings were excluded from the stimulus materials.
We used snippets ( snippet pairs), each pair consisting of one confusing and one clean snippet, which are semantically equivalent and differ only in the presence of an atom of confusion. We presented only one condition of each snippet pair to each participant ( snippets), this way, avoiding learning effects. In total, each participant viewed an equal number of snippets for both conditions ( snippets each).
@@ 4.4 Procedure
At the beginning of the experiment, all participants provided their informed consent and completed a set of questionnaires, including a basic demographic questionnaire [58, 38]. To assess programming experience, we inquired the years and recent frequency of programming activities, as well as all encountered and recently used programming languages. This provides us with a broad estimate of their capabilities and background [59]. Additionally, participants filled out the Edinburgh Handedness Inventory [60] to ensure right-handedness.
@@
The experiment was divided into three consecutive blocks with trials each. Between each block, participants took a break of ten minutes to reduce learning and fatigue effects. We pseudo-randomized the presentation order with the constraints that no more than three snippets of one condition appeared in a row and that, at least, two other snippets appeared between two snippets, containing the same atom of confusion. Each trial started with the presentation of a fixation cross for . Next, the code snippet was presented (see Figure 2). The participants’ task was to determine the output of variable R (as in Result) in each snippet as fast and correctly as possible. Participants had up to to comprehend each snippet, but they could proceed earlier after, at least, by hitting the enter key. Then, we presented four response options: One option was the correct value, two were incorrect values that would follow typical erroneous thought processes associated with the respective atom (only one for two-choice answer options), and one I don’t know option. The last option was intended to prevent participants from guessing. Finally, we asked participants whether the code snippet was easy or difficult to comprehend, also allowing the I don’t know option.
After the experiment, we conducted a semi-structured interview, including questions on the participants’ views about the experiment and each snippet. Finally, we paid the compensation of 12 euros per hour, calculated on a quarter-hourly basis.
@@ 4.5.1 Behavioral data preprocessing
After extracting the behavioral data, we checked whether each participant was generally able to perform the tasks based on their overall answer correctness, which we compared to the estimated correctness based on chance (i.e., due to snippets providing two and snippets providing three answer options). All participants achieved a correctness of, at least, (). Thus, no participants were excluded on the basis of task performance. One trial had to be excluded from the behavioral analysis due to a hardware malfunction, leaving us with trials for the behavioral analysis. We transformed the subjective difficulty rating into a binary variable ( if easy, for hard, I don’t know or no answer), since we can assume that every participant who did not immediately answer the question with ease has experienced, at least, some difficulty.
@@ 4.5.2 Eye-movement recording
We recorded the eye movements using a Tobii Pro Spectrum eye tracker with a sampling rate of and positioned participants using the Tobii Eye Tracker Manager to ensure an optimal head position. Each block started with calibration and validation checking all validation points to ensure high accuracy and precision. If there was a clear offset in the validation points, some points did not receive any data, or the data for a point were scattered across the screen, we repositioned the participant and restarted with calibration and validation.
@@ 4.5.3 Eye-movement preprocessing
The eye-tracking data, specifically the fixations, are the basis for our FRP analysis to reliably detect potential confusion at the correct moment. Missing a fixation or selecting an incorrect fixation would substantially affect the quality of our FRP analysis. Therefore, only trials with high data quality, including a data correction, were considered in our eye-movement analysis.
To this end, we separated the raw eye-tracking data of all trials based on the start and end of the snippet presentations. While the quality was overall high, for some trials, the eye tracker failed to identify the gaze position (i.e., resulting in missing values in the data). Because this poses a risk of missing a relevant fixation in the area of interest, we first excluded trials in which more than of the snippet presentation consisted of missing values. We also removed trials that contained one or more intervals of missing values exceeding one second. Based on this, trials were excluded and trials were left for further analyses.
@@
Another risk to data quality arises from fixations with an offset on the y-axis, because this can lead to assigning fixations to incorrect code lines. To reduce the risk, we used two automated fixation correction algorithms described by Carr et al. [62], which are suitable for code reading (i.e., no assumption of linear line-by-line reading), namely cluster and stretch. We fine-tuned these algorithms, because they both tend to place fixations in as many code lines as possible, although participants may ignore code lines that do not hold information that is relevant for the program execution and variable values (i.e., lines containing only brackets { or }). Specifically, we derived variants that ignore such code lines in favor of other line matches. For each trial, we generated plots to compare the results of the four fixation corrections against the original fixation positioning and, for each fixation correction, to assess the precision on the code lines and suitability for detecting fixations within or near the area of interest later on. Two members of the research team manually reviewed the automated fixation corrections independently of each other. Upon a disagreement ( out of trials), the differences were discussed in detail, on occasions with a third team member to reach consensus, resulting in an almost perfect inter-rater reliability of determined by Cohen’s kappa [63]. During these reviews, we detected that, for a few participants, there was also a noticeable offset on the x-axis. Since this offset was usually constant within a block, we corrected it with a constant offset.
We excluded trials from further analysis when we were unable to confidently align the fixations to the code ( out of ). This was the case when fixations were either entirely off the visualized code and could not be mapped to their correct position, or where single fixations could not be confidently assigned to the correct line. A visualization of all original fixations contrasted with our correction is part of the replication package.
@@ 4.5.4 EEG recording
We recorded continuous EEG from scalp electrodes (Fp1, Fp2, F7, F3, Fz, F4, F8, FC5, FC3, FCz, FC4, FC6, T7, C3, Cz, C4, T8, CP3, CPz, CP4, P7, P3, Pz, P4, P8, O1, O2, and A2) placed in a cap according to the International 10–20 system [64] using BrainVision Recorder (Brain Products). Electrode AFz served as the ground electrode. Additional four electrodes were placed to the left of the left eye as well as above, beneath, and to the right of the right eye monitoring eye movements and blinks (electrooculography; EOG). Two more electrodes were placed over the left and right mastoid. Electrodes were referenced to the electrode placed over the left mastoid during recording. EEG signals were amplified with a bandpass filter of to using a 16 bit BrainAmp DC amplifier. We treated all electrode locations with gel and cotton swabs to achieve an impedance below for EEG and below for EOG electrodes. Continuous analog-to-digital conversion of the EEG and stimulus trigger codes was performed at a sampling frequency of .

### v3

@@ 3 Discussion
The primary goal of this study was to apply an FRP approach – an established technique from psycholinguistics – to achieve a better understanding of the neurocognitive processes that occur when reading confusing program code (i.e., code that contains atoms of confusion) by exploring parallels in terms of ERP components between the processing of program code and natural language processing in an ecologically valid experimental setting. Therefore, we investigated whether atoms of confusion trigger electrophysiological responses, which turned out to be similar to those engaged by unexpected but plausible inputs during language comprehension. In particular, our FRP analysis revealed a late positivity with a frontal scalp topography. It onsets around after the first eye fixation in the area containing the atom of confusion and was significantly larger for confusing code compared to clean code snippets. Notably, this component resembles other late frontal positivities found in natural language processing studies, in which the late frontal positivity is elicited by unexpected words, often but not necessarily in highly constraining contexts [45, 46, 47]. As outlined in Section 1, the late frontal positivity can be dissociated from the P600, a late ERP component, that is elicited by words that are perceived as highly implausible during online comprehension [27, 48].
@@
So what does the late frontal positivity reveal about the processes involved in processing an atom of confusion? In the literature on natural language comprehension, the predominant account for a late frontal positivity holds that it reflects expectancy mismatch related processing [45, 46] or late processing costs due to the perceived difference between a prediction and the bottom-up input in highly constraining sentences [49, 50]. However, a range of studies has also observed that unexpected words within weakly constraining sentences can elicit a late frontal positivity [51, 52, 26, 53, 54, 55]. Brothers et al. [34, 56] proposed that the essential properties of a stimulus that lead to a late frontal positivity might be the presence of an unexpected but plausible word, which causes a substantial update of the situation model. More specifically, Kuperberg et al. [57] suggest that late frontal positivities arise in situations where there is initial uncertainty about the situation model, which triggers the retrieval of information from long term memory, enabling the comprehender to accumulate confidence in their ability to successfully update their situation model. In the case of high contextual constraint and unexpected lexical material, this occurs because the unexpected word first reduces confidence in the situation model, triggers retrieval from long-term memory, and finally leads to a high confidence decision that the situation model can be successfully updated [57]. In the case of a low-constraint context and unexpected but plausible lexical input, retrieval of relevant event-level information from long-term memory allows the comprehender to accumulate confidence about how to update their situation model. Kuperberg et al. [57] point out that the late frontal positivity effect in language processing seems to depend on the task, how engaged an individual is in deeply understanding the situation, and whether they can achieve confidence about the situation model based on the information they retrieved from long-term memory.
@@
These ideas align well with a plausible mechanism for processing the atoms of confusion: Atoms of confusion are infrequent in code [11] and hence generally unexpected, but they are neither implausible nor incorrect, as they reflect syntactically correct code patterns that can be unambiguously executed by the computer. However, when a comprehender encounters an atom of confusion, they are likely to be unsure at first about its effect on program execution and the variables they need to keep track of for the task (i.e., they are unsure about how to update their situation model), which is in line with observations that relate confusing code patterns to increased cognitive load [16]. As a result, comprehenders access their long-term memory to remember exactly how to interpret the atom of confusion. If this retrieval is successful, the comprehender can confidently update their situation model. In the case of code without an atom, the situation model will also be updated, but there is no point of uncertainty followed by confidence on how to do that. This account also generates new predictions to be tested in future studies: We would expect that the late frontal positivity should only be present in trials where the comprehender could successfully resolve the ambiguity. It should therefore not appear in situations in which the atom of confusion contains a syntactically invalid code element, for which we would expect a P600 with a posterior scalp topography [42]. Also, we do not expect a late frontal positivity to appear when the comprehender remains uncertain about the semantics of the atom or when they are given a different task that does not require updating the situation model.
@@
In summary, we presented, to our knowledge, the first-ever study that employs an FRP approach to explore the processing of confusing elements in program code. The results suggest that the brain engages highly similar mechanisms for processing unexpected but plausible elements in program code and natural language comprehension. This implies that studies on program comprehension can rely on established ERP components in natural language processing. The FRP approach offers new research perspectives with high ecological validity, focusing on the temporal characteristics of online program code comprehension. These results are significant for software engineering research, as they provide a foundation for a better understanding of code comprehension. Based on this common methodological ground, methods and findings from program comprehension research may be transferred back to research on natural language comprehension. The findings of our study, in particular, and this line of research, in general, also have far-reaching implications for the practice of software engineering. By shedding light on how programmers’ brains process and respond to confusing code, we can enable programmers to develop and maintain more secure and reliable software. Specifically, our line of research supports programming practice by providing empirical evidence for coding conventions that avoid confusion-inducing code patterns. Moreover, programming language designers can leverage these insights to design consistent language constructs that minimize the potential for confusion and resulting severe mistakes. Finally, educators can incorporate these findings into their teaching practices, ultimately educating future programmers to apply more effective software development practices.
@@ 4.1 Participants
To estimate the number of participants required to find the hypothesized effect, we performed a power analysis with , , and (based on the reported ) for the late frontal positivity detected by Thornhill and Peterson [26] for high-constraint unrelated sentence endings. This analysis estimated a minimum of participants to correctly measure an existing effect (two-sided t-test). Due to various uncertainties in our setup and for the counterbalancing of the three version blocks across participants, we increased our sample to participants. While this number might appear low, it is in line with ERP research in general [58, 59, 46, 44] and specifically with existing studies in the domain of software engineering [60, 61]. Further studies using larger samples shall be performed to generalize our results to a larger population, with an explicit focus on expert developers.
We recruited participants at our host institution (Saarland University) via e-mail lists, flyers, and (online) bulletin boards. We required participants to have at least an undergraduate degree in computer science or a related field. We further asked for basic experience with an object-oriented programming language (e.g., Java or C++), as the selected snippets use minimal language-specific operations. We only invited right-handed participants.
@@
Overall, we invited participants to our experiment. To ensure high data quality, we excluded two participants from our analysis due to hardware malfunctions during the experiment or frequent excessive artifacts in the EEG data (see Section 4.5.4 and 4.5.5). out of the included participants were female, and were male, with a total age range from to years, and participants being younger than . In Table 2, we provide an overview of our included participants’ demographics and their programming experience.
@@ 4.2 Experimental variables
The independent variable of our experiment is the snippet type, which has two conditions: either confusing or clean. As dependent variables, we measured the comprehension time, answer correctness, subjective difficulty rating, eye movements via an eye tracker, and brain activity via an EEG system for each snippet (see Figure 2).
@@ 4.3 Stimuli
The code snippets in our study are based on the snippet pairs from the study by Langhout and Aniche [13], which we adjusted to reduce fixations and confounding neurocognitive processes in the FRP analysis. In a pilot study, we rated these snippet pairs considering three criteria regarding the snippet structure, the area of the atom of confusion, and behavioral response (cf. supplementary information, Appendix A). Specifically, we required that within each snippet pair, the snippets have a comparable structure to reduce the influence of confounding factors. This was determined by the number and length of the code lines. Furthermore, to facilitate the selection of suitable fixations for FRP analysis, each snippet in a pair must contain a single clear area of interest that either includes the atom of confusion (for confusing snippets) or a clarified equivalent (for clean snippets). Finally, we tested whether the snippets were of similar complexity and whether they were, in principle, understandable to students of our target population (familiarity with the syntax of Java) in a pilot study. To this end, we set a threshold of answer correctness or subjective difficulty rating as easy for either condition. Snippets that showed a clear violation of the expected behavior by lower correctness or fewer easy ratings were excluded from the stimulus materials.
We used snippets ( snippet pairs), each pair consisting of one confusing and one clean snippet, which are semantically equivalent and differ only in the presence of an atom of confusion. We presented only one condition of each snippet pair to each participant ( snippets), thereby avoiding learning effects. In total, each participant viewed an equal number of snippets for both conditions ( snippets each).
While the combined eye-tracking and EEG approach offers a higher ecological validity than using a token-wise presentation which is common in ERP studies with linguistic events, it is important to acknowledge that constraints inherent to eye tracking and EEG remain. These constraints limit the length and, consequently, the ecological validity of the presented code snippets. Nevertheless, because atoms of confusion consist of small, localized patterns, they can be meaningfully embedded within code even under these length limitations.
@@ 4.4 Procedure
At the beginning of the experiment, all participants provided their informed consent and completed a set of questionnaires, including a basic demographic questionnaire [62, 38]. To assess programming experience, we inquired about the years and recent frequency of programming activities, as well as all programming languages encountered and recently used. This provides us with a broad estimate of their capabilities and background [63]. Additionally, participants filled out the Edinburgh Handedness Inventory [64] to ensure right-handedness.
@@
The experiment was divided into three consecutive blocks with trials each. Between each block, participants took a break of ten minutes to reduce learning and fatigue effects. We pseudo-randomized the presentation order with the constraints that no more than three snippets of one condition appeared in a row and that at least two other snippets appeared between two snippets containing the same atom of confusion. Each trial started with the presentation of a fixation cross for . Next, the code snippet was presented (see Figure 2). The participants’ task was to determine the output of variable R (as in Result) in each snippet as quickly and accurately as possible. Participants had up to to comprehend each snippet, but they could proceed earlier after at least by hitting the enter key. Then, we presented four response options: One option was the correct value, two were incorrect values that would follow typical erroneous thought processes associated with the respective atom (only one for two-choice answer options), and one I don’t know option. The last option was intended to prevent participants from guessing. Finally, we asked participants whether the code snippet was easy or difficult to comprehend, also allowing the I don’t know option.
After the experiment, we conducted a semi-structured interview that included questions on the participants’ views about the experiment and each snippet. Finally, we paid the compensation of 12 euros per hour, calculated on a quarter-hourly basis.
@@ 4.5.1 Behavioral data preprocessing
After extracting the behavioral data, we checked whether each participant was generally able to perform the tasks based on their overall answer correctness, which we compared to the estimated correctness based on chance (i.e., due to snippets providing two and snippets providing three answer options). All participants achieved a correctness of at least (). Thus, no participants were excluded on the basis of task performance. One trial had to be excluded from the behavioral analysis due to a hardware malfunction, leaving us with trials for the behavioral analysis. We transformed the subjective difficulty rating into a binary variable ( if easy, for hard, I don’t know , or no answer), since we can assume that every participant who did not immediately answer the question with ease has experienced at least some difficulty.
@@ 4.5.2 Eye-movement recording
We recorded the eye movements using a Tobii Pro Spectrum eye tracker with a sampling rate of and positioned participants using the Tobii Eye Tracker Manager to ensure an optimal head position. Each block began with calibration and validation, checking all validation points to ensure high accuracy and precision. If there was a clear offset in the validation points, if some points did not receive any data, or if the data for a point were scattered across the screen, we repositioned the participant and restarted with calibration and validation.
@@ 4.5.3 Eye-movement preprocessing
The eye-tracking data, specifically the fixations, are the basis for our FRP analysis to reliably detect potential confusion at the correct moment. Missing a fixation or selecting an incorrect fixation would substantially affect the quality of our FRP analysis. Therefore, only trials with high data quality, including data correction, were considered in our eye-movement analysis.
To this end, we separated the raw eye-tracking data of all trials based on the start and end of the snippet presentations. While the quality was overall high, for some trials, the eye tracker failed to identify the gaze position (i.e., resulting in missing values in the data). Because this poses a risk of missing a relevant fixation in the area of interest, we first excluded trials in which more than of the snippet presentation consisted of missing values. We also removed trials that contained one or more intervals of missing values exceeding one second. Based on this, trials were excluded, and trials were left for further analyses.
@@
Another risk to data quality arises from fixations with an offset on the y-axis, because this can lead to assigning fixations to incorrect code lines. To reduce the risk, we used two automated fixation correction algorithms described by Carr et al. [66], which are suitable for code reading (i.e., no assumption of linear line-by-line reading), namely cluster and stretch. We fine-tuned these algorithms, because they both tend to place fixations in as many code lines as possible, although participants may ignore code lines that do not hold information relevant for program execution and variable values (i.e., lines containing only brackets { or }). Specifically, we derived variants that ignore such code lines in favor of other line matches. For each trial, we generated plots to compare the results of the four fixation corrections against the original fixation positioning and, for each fixation correction, to assess the precision on the code lines and suitability for detecting fixations within or near the area of interest later on. Two members of the research team manually reviewed the automated fixation corrections independently of each other. Upon a disagreement ( out of trials), the differences were discussed in detail, on occasions with a third team member to reach consensus, resulting in an almost perfect inter-rater reliability of determined by Cohen’s kappa [67]. During these reviews, we detected that, for a few participants, there was also a noticeable offset on the x-axis. Since this offset was usually constant within a block, we corrected it with a constant offset.
We excluded trials from further analysis when we were unable to confidently align the fixations to the code ( out of ). This was the case when fixations were either entirely off the visualized code and could not be mapped to their correct position, or when single fixations could not be confidently assigned to the correct line. A visualization of all original fixations contrasted with our correction is part of the replication package.
@@ 4.5.4 EEG recording
We recorded continuous EEG from scalp electrodes (Fp1, Fp2, F7, F3, Fz, F4, F8, FC5, FC3, FCz, FC4, FC6, T7, C3, Cz, C4, T8, CP3, CPz, CP4, P7, P3, Pz, P4, P8, O1, O2, and A2) placed in a cap according to the International 10–20 system [68] using BrainVision Recorder (Brain Products). Electrode AFz served as the ground electrode. An additional four electrodes were placed to the left of the left eye as well as above, beneath, and to the right of the right eye to monitor eye movements and blinks (electrooculography; EOG). Two more electrodes were placed over the left and right mastoids. Electrodes were referenced to the electrode placed over the left mastoid during recording. EEG signals were amplified with a bandpass filter of to using a 16 bit BrainAmp DC amplifier. We treated all electrode locations with gel and cotton swabs to achieve an impedance below for EEG and below for EOG electrodes. Continuous analog-to-digital conversion of the EEG and stimulus trigger codes was performed at a sampling frequency of .